QUESTION 1

Use the method of First Principles to find f'(x), given that

f(x) = x^2 + 2x
1
Expert's answer
2019-08-12T09:25:20-0400

"f'(x)=lim_{h\to 0}\frac{(x+h)^2+2(x+h)-x^2-2x}{h}=lim_{h\to 0}\frac{2xh+2h+h^2}{h}=lim_{h\to 0}(2x+2+h)=2x+2"